Timberly Construction negotiates a lump-sum purchase of several assets from a company that is going out of business. The purcApportioned Cost Allocation of Total Appraised Value Percent of Total Total cost of х Acquisition Cost Appraised Value % X BuRecord the costs of lump-sum purchase. Note: Enter debits before credits. Date General Journal Debit Credit Jan 01Compute the depreciation expense for year 2017 on the building using the straight-line method, assuming a 15-year life a a $2Compute the depreciation expense for year 2017 on the land improvements assuming a five-year life and double-declining balanc
Answer #1

1) Allocation

Allocation of total Cost Appraised value Percent of total appraised value Total Cost of acquisition Apportioned Cost
Building 508800 53% 900000 477000
Land 297600 31% 900000 279000
Land improvement 28800 3% 900000 27000
Vehicle 124800 13% 900000 117000
Total 960000 100% 900000

2) Journal entry

date account and explanation Debit Credit
Jan 1 Building 477000
Land 279000
Land improvement 27000
Vehicle 117000
Cash 900000

3) Depreciation expense on building = (477000-27000/15) = 30000

4) Depreciation expense on land improvement = 27000*40% = 10800